OpenStudy (lonjay24):

The mass of a sample of sodium bicarbonate is 2.1 kilograms (kg). There are 1,000 grams (g) in 1 kg, and 1 mc020-1.jpg 109 nanograms (ng) in 1 g. What is the mass of this sample in ng? 2.1 mc020-2.jpg 103 ng 2.1 mc020-3.jpg 106 ng 2.1 mc020-4.jpg 109 ng 2.1 mc020-5.jpg 1012 ng

OpenStudy (kkutie7):

stoichiometry. I don't know what the deal with your mc20-jpgs but I'll still help you. I'ts important to get a visual.

OpenStudy (kkutie7):

set things up like fractions. \[\frac{2.1kg}{1}*\frac{1,000g}{1kg}*\frac{109ng}{g}\] the way I have set it up makes it easy to see that anything in denominator can cancel out things in the numerator. By writing the units like this you can cross things out until you are left with the units you are looking for: \[\frac{2.1}{1}*\frac{1,000}{1}*\frac{109ng}{1}\] now all you have to do is multiply.